Can someone help me solve this problem?
I need to solve the following equation for x, and show proof:
7(3x + 4) = 8(2x + 5) + 13
Thanks,
Sher

Answer by stanbon(75887)

7(3x + 4) = 8(2x + 5) + 13
21x+28 = 16x+40+13
21x+28 = 16x+53
5x=25
x=5
Prove It:
7(3(5)+4) = 8(2(5)+5) + 13
7(15+4) = 8(10+5) + 13
7(19) = 8(15)+13
133 = 133
Cheers,
Stan H.